7, 16, 8, 27, 9, ?
Answer by farohw(175)


7, 16, 8, 27, 9, (40), (10)...

It appears that every other number starting at 7 is consecutive. Hence, 7, 8, 9, 10. However 16, 27, 40 are multiplied by 2*8 = 16, 3*9 = 27 and 4*10 = 40. Many times these things are solved through trial and error, sometimes you have to work backwards so to speak but do you see the pattern?

Best,
Farohw